Solve cos2x+sinx=1in[0,2pi)?
Answer by stanbon(75887)

Solve cos(2x)+sinx=1in[0,2pi)?
-----
cos(2x) = cos^2(x)-sin^2(x) = 1-2sin^2(x)
--------------------------
Your equation:
1-2sin^2(x)+sin(x) = 1
-----
Rearrange:
2sin^2(x)-sin(x) = 0
----
Factor:
sin(x)[2sin(x)-1] = 0
-----
sin(x) = 0 or sin(x) = 1/2
-------
x = 0 or pi or pi/6 or (5/6)pi
=====================================
Cheers,
Stan H.
-----------------------
------